Use the bisection to find the positive root of the equation to the nearest tenth.
x^2-11=0 (Show procedures)

Answer by stanbon(75887)

Use the bisection to find the positive root of the equation to the nearest tenth.
x^2-11=0 (Show procedures)
----
Factor::
(x-sqrt(11))(x+sqrt(11)) = 0
----
x = sqrt(11) or x = -sqrt(11)
-----
3^2 < 11 < 4^2
------
a------b-----(a+b)/2---[(a+b)/2]^2
3------4------3.5-------3.5^2 = 12.25---- > 11
3-----3.5-----3.25------3.25^2 = 10.56---- < 11
3.25--3.5-----3.375-----3.375^2 = 11.39---- > 11
etc
----
sqrt(11) = 3.32
=======================
-sqrt(11) = -3.32
=======================
Cheers,
Stan H.
------------